OpenStudy (anonymous):

Suppose that y varies directly with x and inversely with z, and y = 18 when x = 15 and z = 5. Write the equation that models the relationship. Then find y when x = 21 and z = 7.

hartnn (hartnn):

y directly varies with x, so u can write y=mx y inversely varies with z, so u can write y=n/z overall u can write \(y=\frac{kx}{z}\) now put,y=18,x=15,z=5 to find k. can u find k ?

OpenStudy (anonymous):

So Far i have 18=k5/k

OpenStudy (anonymous):

were do i go from here

hartnn (hartnn):

how did u get that? 18=k(15/5) 18=k(3) k=18/3=6 ok till here?